# Vision-Language Models for Edge Networks: A Comprehensive Survey

### v1

@@ V-B Visual-Language Models Optimization for Edge Networks
How do we optimize VLMs models architecture for edge devices in terms of performance and complexity? Another technical trend involves optimizing the visual encoder and language model components. Researchers are focusing on balancing these two components to achieve efficiency in resource-constrained environments. For example, the Imp project explores the use of smaller LLMs like Phi-2 and optimized visual encoders like SigLIP, which perform better than traditional CLIP-based encoders [156, 168]. This results in a better generalization when deployed on edge devices. These advancements significantly reduce the required computational power, making models more suitable for mobile and embedded systems. Many recent applications of lightweight VLMs are tailored for specific edge use cases, such as real-time analysis for drones, robots, and surveillance systems. For instance, the Moondream2 model is designed to be highly efficient and able to process complex vision-language tasks like interpreting security footage and performing remote inspections [157, 155]. These models run on as little as $5$GB of memory, making them ideal for fully remote use cases where continuous connectivity cannot be guaranteed. A significant technical trend is the adaptation of VLMs to work with various modalities, such as depth and thermal cameras, alongside traditional RGB images. This cross-modality adaptation is crucial for applications in autonomous systems, where VLMs must process diverse visual inputs. EdgeVL, for instance, employs cross-modality learning, enabling efficient operation across different input types while preserving performance, which is essential for robots and autonomous systems operating in dynamic environments [156, 169].
@@ V-C Distributed Implementation of Edge VLMs
How do we enable distributed implementation of VLMs for various applications? Federated learning and edge computing are emerging as critical enablers for distributed lightweight VLMs. Federated learning allows models to be fine-tuned directly on edge devices without transferring sensitive data to the cloud, preserving user privacy and reducing latency [156]. This is particularly important in healthcare and security applications with high data sensitivity. In parallel, edge computing enables these models to process data closer to the source, improving response times and making real-time decision-making more feasible [157]. Besides the applications we mentioned earlier, there are some trends in which VLMs are gaining significant traction due to their ability to perform complex tasks on resource-constrained edge devices. In autonomous systems such as drones, robots, and autonomous vehicles, VLMs are utilized for real-time object detection, scene understanding, and navigation, allowing these systems to operate without reliance on cloud computing. For example, drones equipped with VLMs can monitor wildlife, inspect infrastructure, and assess environmental conditions, which has crucial implications for disaster relief and agriculture [156].
@@ V-D Context-Aware VLMs for Edge Networks
How do we propose VLMs for edge networks with context-awareness? Extending VLMs to the network edge by using the edge data for further training (i.e., context-aware) of pre-trained models is necessary for many applications to adapt to specific scenarios. Meanwhile, the models deployed at the network edge must be lightweight during extension. Recent advancements in context-aware VLMs have emphasized the development of lightweight, task-specific architectures suited for edge networks where computational resources are limited. These models, such as EM-VLM4AD, MiniDrive, and LiteViLA, incorporate several techniques to efficiently manage multimodal data by reducing model complexity without significantly compromising performance [170, 171, 172]. One of the core techniques used in lightweight VLMs involves leveraging efficient image embedding mechanisms, such as ViT-based patch projections and gated pooling attention, allowing multi-view image data to be processed with minimal latency. EM-VLM4AD, for example, flattens image patches and performs gated pooling to facilitate processes requiring fewer resources to produce a single representation by compressing and summarizing multiple views before fusing a language model (e.g., T5-base) for question-answering based tasks in autonomous driving applications [170]. It reduces inference time and achieves superior accuracy in several tasks, especially in path planning and traffic behavior analysis. For example, LiteViLA employs a Mixture of Adapters (MoA) approach, dynamically activating lightweight adapters specialized for individual subtasks, including object detection and scene understanding, resulting in efficient resource allocation and allowing for a diverse array of tasks to be performed [171]. LiteViLA supports different edge tasks in autonomous systems, and such modularity provides robustness for every operational condition. Moreover, models originally proposed for drones [172] also combine lightweight portions, such as YOLOv7 detectors, with VLM architectures to generate real-time object detection and scene description. These models focus on low latency by implementing simple structures of encoder-decoder and use quantized or pruned LLMs (GPT-3, TinyLLaVA) for the language, which makes them ideal for use in applications where power and performance are the most significant constraints.
@@ V-E Cross-Modality Learning and Adaptation for Multi-Sensor Applications
How does one enable VLMs to effectively integrate and adapt to diverse sensor modalities, such as thermal, depth, and hyperspectral data in many edge applications? Cross-modality learning and adaptation have emerged as one of the most recent and important technologies in promoting the advancement of VLMs to work across various sensor types, including thermal, depth, and hyperspectral modalities. It uses data observed from multiple domains to boost the ability of VLMs to perceive a scene, which is a fundamental requirement for tasks related to autonomous systems, robot control, and environmental monitoring. For example, the models ViPT [173], UC2 [174], and CMT [175] employed cross-modal fusion methods combining thermal and RGB data into a shared feature space that enables coherent interpretation in all respective modalities. ViPT is designed to build on pre-trained RGB-based models that can serve as multi-modal backbones for new tasks like tracking, where RGB and depth data can be fused through a transformer-based encoder that incorporates modality-complementary prompters [174]. Likewise, depth estimation models commonly rely on RGB and thermal data fusion using dedicated networks, such as a 3D cross-modal transformation module, which aligns data from separate modalities to increase depth prediction accuracy in dim lighting scenarios [175]. These models have recently been explored to generate confidence maps and align the modalities originating from the various sensors to select the most accurate one, thus enhancing outputs from the complex multi-sensor setups. These cross-modal approaches have proven successful in various fields outside of robotics. Analogous fusion techniques in environmental monitoring and agriculture would allow models to deal with multimodal data, for example, from drones (hyperspectral imagery) or IoT sensors instead of drone imagery more tightly connected to the ecosystem monitoring and precision agriculture use cases [173, 172].
@@ V-F Security
How do we enable edge VLMs while ensuring security? The deployment of lightweight VLMs in both cloud and edge environments introduces significant security challenges, as these models are susceptible to a range of attacks. Due to their large-scale usage, VLMs are exposed to risks like model inversion attacks, which allow adversaries to reconstruct training data from model outputs, compromising privacy. This vulnerability arises from their reliance on shared representations across modalities, making them targets for privacy-related attacks [176, 177]. Recent research has focused on enhancing robustness against adversarial attacks. Adversarial training incorporates noise into predictions, reducing the effectiveness of adversarial examples. Robust gradient masking methods limit adversaries’ ability to exploit gradients [178, 177]. Ensemble-based defenses, such as randomized input transformations and multi-layer protection, provide a layered security approach. Additionally, using adversarial examples to ”boost” defenses prevents unauthorized data reconstruction from VLM outputs [176, 178].
When VLMs are deployed in distributed settings, ensuring secure data transmission between edge devices and cloud servers is critical. Without proper encryption, attackers can intercept or manipulate transmitted data. Secure communication protocols like AES-256 encryption and lightweight frameworks ensure robust protection without performance loss [179]. Dynamic key exchange protocols provide real-time key generation, improving resistance to attacks. Secure Multiparty Computation (SMC) techniques allow joint computations without revealing inputs [180]. Edge VLMs are prone to hardware attacks, including tampering and malware. Trusted Execution Environments (TEEs) like ARM TrustZone mitigate these risks by securely running critical components, even in untrusted environments [177]. Blockchain-based solutions enhance protection against unauthorized firmware updates, maintaining edge device integrity [181].
Poisoning attacks involve inserting malicious data into VLM training sets, degrading performance. Federated learning environments are especially vulnerable. Byzantine-resilient federated systems detect and eliminate poisoned data without affecting performance [182]. Anomaly detection methods effectively filter out malicious updates. Securing aggregation of model updates in federated learning is vital to prevent data inference or performance degradation. Differential privacy techniques obscure individual updates while maintaining accuracy. Homomorphic encryption protects data during model aggregation [183]. Blockchain-based systems ensure transparency and prevent tampering [180]. Trust frameworks and blockchain-based governance models maintain transparency in access control and secure deployment of VLMs. Ethical governance frameworks ensure compliance with security standards [181].
@@ V-G Privacy
How do we propose privacy-aware VLMs? The rapid adoption of VLMs in cloud-based systems has heightened concerns about user privacy. VLMs process multimodal data—such as personal images and text—often transmitted to remote servers, raising challenges in ensuring data confidentiality. Safeguarding sensitive information against unauthorized access is critical. Recent advancements focus on privacy-preserving techniques that enable secure inference and data handling without compromising performance. Federated learning improves privacy by keeping data localized while sharing only model updates, but privacy concerns remain due to possible inference of sensitive information from these updates. [184] proposed a homomorphic encryption-based framework, allowing model updates to be computed without exposing raw data. Earlier, [185] introduced the PFMLP framework using partially homomorphic encryption to protect data during federated learning with minimal accuracy loss. In 2023, [186] introduced a privacy-preserving inference framework combining homomorphic encryption and random privacy masks, preventing access to raw input data with low computational overhead. Similarly, [187] integrated secure multiparty computation (SMC) with homomorphic encryption to enhance federated learning systems.
Another privacy issue in cloud-based VLMs is the risk of membership inference attacks, where VLMs trained on private data collections lead to privacy concerns [188]. [188] introduced a federated learning framework encrypting model updates with homomorphic encryption to prevent sensitive data leakage. [189] proposed a multi-key encryption design protecting against membership inference attacks by ensuring updates are encrypted and inaccessible to a single participant. Additionally, cloud-based VLMs need to address data ownership and control. Once uploaded, users lose control over their data. [190] developed a federated learning framework for IoT systems, decentralizing client data and safeguarding against collusion attacks. Privacy laws like GDPR impose further constraints. [191] proposed a framework combining differential privacy with homomorphic encryption for GDPR compliance, maintaining data utility without sacrificing model performance.
@@ V-H Communication Model for Edge VLMs
How does one enable communication resources efficient edge VLMs? Training requires significant communication resources, especially for distributed VLMs at the network edge. Offloading model processing to the cloud introduces communication overhead. [192] proposed an optimized distributed CNN framework to reduce memory footprint and communication overhead in edge-cloud setups. [193] introduced a declarative framework optimizing data flows between edge devices and the cloud. Energy consumption is another major challenge. In 2023, [194] introduced an energy-efficient framework for NLP on edge devices leveraging heterogeneous memory architectures to reduce energy consumption while maintaining high performance. [195] developed EdgePipe, a distributed framework using pipeline parallelism to improve energy efficiency during inference, achieving speedups without sacrificing accuracy.
Real-time inference is critical for edge applications. [196] proposed DeViT, a framework decomposing large vision transformers into smaller models for collaborative inference on edge devices, reducing latency and communication overhead. [195] also showed pipeline parallelism could speed up inference on heterogeneous edge devices, achieving high throughput with negligible accuracy drop. Generalizing VLMs across heterogeneous devices presents another challenge. [197] proposed DCA-NAS, enabling neural architecture search for diverse hardware configurations, allowing fine-tuned model designs for varying constraints. [195] demonstrated pipeline parallelism’s adaptability to heterogeneous hardware, improving performance and flexibility.

### v2

@@ V-B Visual-Language Models Optimization for Edge Networks
How do we optimize VLMs models architecture for edge devices in terms of performance and complexity? A key trend involves optimizing the visual encoder and language model components to balance performance and efficiency in constrained environments. For instance, the Imp project uses smaller LLMs like Phi-2 and optimized visual encoders such as SigLIP, which outperform traditional CLIP-based models [155, 166], enabling better generalization on edge devices. These improvements significantly reduce computational requirements, making the models suitable for mobile and embedded systems.
Lightweight VLMs are increasingly tailored for edge-specific use cases such as real-time analysis for drones, robots, and surveillance. The Moondream2 model, for example, is designed for efficiency in complex tasks like interpreting security footage and conducting remote inspections [156, 154]. It operates with as little as $5$GB of memory, making it ideal for remote settings with limited connectivity. Another important direction is enabling VLMs to handle multiple modalities, including depth and thermal imaging in addition to RGB. This cross-modality adaptation is critical for autonomous systems that must process diverse visual inputs. The EdgeVL framework applies cross-modality learning to maintain performance across input types, supporting robust operation in dynamic edge environments like robotics and autonomous navigation [155, 167].
Problems in Model Optimization for Edge Networks Deploying advanced optimization techniques such as dynamic inference and model scaling in edge networks introduces multiple practical challenges. Dynamic inference, which adjusts model complexity at runtime based on available resources, faces critical issues, including limited computational resources and unpredictable latency, adversely impacting real-time performance [168]. These can be mitigated through modular architectures that selectively activate model components depending on resource availability and by employing resource-aware scheduling algorithms for effective real-time adjustments [168]. Similarly, model scaling—tailoring models to match diverse edge hardware capabilities—is complicated by hardware heterogeneity and significant maintenance overhead due to managing multiple model variants [168, 169]. This challenge can be addressed by efficiently employing automated model compression techniques, like pruning and quantization, to create device-specific models [169]. Additionally, federated learning methods enable collaborative training across edge devices, effectively adapting models to varied hardware configurations without needing centralized data collection [168].
@@ V-C Distributed Implementation of Edge VLMs
How do we enable distributed implementation of VLMs for various applications?. Federated learning and edge computing are emerging as critical enablers for distributed lightweight VLMs. Federated learning allows models to be fine-tuned directly on edge devices without transferring sensitive data to the cloud, preserving user privacy and reducing latency [155]. This is particularly important in healthcare and security applications with high data sensitivity. In parallel, edge computing enables models to process data closer to the source, improving response times and enabling real-time decision-making [156]. Beyond previously mentioned applications, VLMs are gaining traction in resource-constrained edge environments due to their ability to perform complex tasks. In autonomous systems like drones, robots, and vehicles, VLMs support real-time object detection, scene understanding, and navigation—allowing operation without cloud dependence. For example, drones equipped with VLMs can monitor wildlife, inspect infrastructure, and assess environmental conditions—crucial for disaster relief and agriculture [155].
Integration of Federated Learning and VLMs
Integrating FL with VLMs enables privacy-preserving and resource-efficient intelligence in edge computing. Recent research explores adapting VLMs to federated settings:
Federated Prompt Learning: Instead of fine-tuning entire VLMs, this approach optimizes only prompt parameters, reducing communication and computational overhead, making deployment feasible on resource-constrained devices [170].
Personalized Federated Learning: Data heterogeneity across clients challenges global model generalization. Techniques like Federated Mixture of Experts enable personalized model adaptations by learning tailored expert combinations [171].
Parameter-Efficient Fine-Tuning: Methods such as LoRA facilitate adapting VLMs in FL without updating the entire model, reducing memory usage and speedup training [172].
Communication Efficiency: The large size of VLMs increases communication costs. Techniques like gradient compression, selective updates, and client sampling mitigate these issues while maintaining performance [173].
Multimodal Data Handling: Clients may have unpaired or uni-modal data, complicating multimodal VLM training. Approaches like Federated Modality Complementary and Collaboration (FedCola) align and integrate different modalities across clients [174].
Despite these advancements, challenges persist in FL-VLM integration. Data heterogeneity can lead to model divergence, requiring robust aggregation and personalization techniques. Limited computational resources at the edge necessitate lightweight architectures and efficient training strategies. Ensuring privacy remains critical, with differential privacy and secure aggregation methods improving data protection.
Future research should focus on scalable federated fine-tuning, cross-client multimodal alignment, and optimizing FL for real-time edge applications.
@@ V-D Context-Aware VLMs for Edge Networks
How do we propose VLMs for edge networks with context-awareness? Extending VLMs to the edge using local data for context-aware training is essential for adapting to specific scenarios. These models must remain lightweight during deployment and extension. Recent advances emphasize task-specific, low-complexity architectures suited for resource-constrained edge environments. Models like EM-VLM4AD, MiniDrive, and LiteViLA adopt strategies to handle multimodal data efficiently while preserving performance [175, 176, 177]. A core technique is efficient image embeddings, such as ViT-based patch projections and gated pooling attention, which enables fast processing of multi-view images. EM-VLM4AD flattens image patches and applies gated pooling to produce a compressed representation, later fused with a language model for tasks like question answering in autonomous driving [175]. This design reduces inference time while improving accuracy in path planning and traffic behavior analysis tasks. LiteViLA uses a Mixture of Adapters (MoA) strategy, activating lightweight adapters tailored to subtasks such as object detection and scene understanding [176]. This modularity ensures efficient resource use and robust performance across diverse edge conditions. For drone-based applications, models combine lightweight components like YOLOv7 with VLMs for real-time object detection and scene description [177]. These use simple encoder-decoder structures and quantized or pruned LLMs (e.g.TinyLLaVA), optimizing for low latency and energy efficiency—ideal for power-constrained environments.
@@ V-E Cross-Modality Learning and Adaptation for Multi-Sensor Applications
How does one enable VLMs to effectively integrate and adapt to diverse sensor modalities, such as thermal, depth, and hyperspectral data in many edge applications? Cross-modality learning and adaptation have become key technologies in enabling VLMs to operate across diverse sensor types, including thermal, depth, and hyperspectral data. By using information from multiple domains, these methods enhance scene understanding—critical for autonomous systems, robotic control, and environmental monitoring.
Challenges in edge environments include: Data Synchronization: Ensuring temporal alignment of data from different modalities is complex, especially when each modality may have varying data rates and latencies [178]. Potential Solutions: Developing lightweight fusion techniques, such as efficient early or late fusion algorithms, can mitigate resource constraints by minimizing computational overhead [179]. Additionally, employing edge-cloud collaboration by offloading certain fusion tasks to the cloud, while retaining latency-sensitive operations at the edge, effectively balances computational load and maintains performance [180]. Models such as ViPT [181], UC2 [182], and CMT [183] employ fusion techniques to integrate thermal and RGB data into shared feature spaces, enabling consistent interpretation across modalities. ViPT builds on pre-trained RGB models and fuses RGB and depth inputs via a transformer encoder with modality-complementary prompters [182]. Similarly, depth estimation models use RGB-thermal fusion, often with a 3D cross-modal transformation module, to improve performance in low-light conditions [183]. These models can generate confidence maps and align sensor modalities to prioritize the most accurate inputs, enhancing outcomes in multi-sensor settings. Cross-modal techniques have also shown promise beyond robotics. In fields like environmental monitoring and agriculture, similar fusion approaches allow models to work with multimodal inputs from drones (e.g., hyperspectral imagery) or IoT sensors, supporting use cases such as ecosystem monitoring and precision agriculture [181, 177].
@@ V-F Security
How do we enable edge VLMs while ensuring security? Deploying lightweight VLMs in cloud and edge environments introduces security challenges, as these models are vulnerable to various attacks. Model inversion attacks, for example, can reconstruct training data from outputs due to shared cross-modal representations [184, 185]. To counter this, recent work uses adversarial training, which introduces noise to reduce attack effectiveness, and gradient masking to limit gradient exploitation [186, 185]. Ensemble-based defenses like randomized input transformations and multi-layer protection further strengthen security. Adversarial examples are also used to prevent unauthorized data reconstruction [184, 186].
In distributed settings, secure data transmission is essential. Without encryption, data can be intercepted or altered. Protocols such as AES-256 and lightweight encryption frameworks offer strong protection with minimal overhead [187]. Dynamic key exchange enhances resistance, while Secure Multiparty Computation (SMC) enables joint computation without revealing inputs [120]. Edge VLMs also face hardware threats like tampering and malware. Trusted Execution Environments (TEEs), such as ARM TrustZone, protect critical components even in untrusted settings [185]. Blockchain-based mechanisms help prevent unauthorized firmware updates and maintain device integrity [188].
Poisoning attacks, where malicious data corrupts training, concern federated learning significantly. Byzantine-resilient systems and anomaly detection methods can identify and remove such inputs [189]. To secure aggregation, differential privacy hides individual updates while preserving accuracy, and homomorphic encryption protects data during computation [190]. Blockchain ensures tamper resistance and transparency [120]. Governance frameworks based on trust and blockchain support secure access control and deployment, while ethical guidelines ensure compliance [188].
@@ V-G Privacy
How do we propose privacy-aware VLMs? The growing adoption of VLMs in cloud-based systems raises serious privacy concerns. These models process multimodal data—such as personal images and text—often transmitted to remote servers, making data confidentiality a critical issue. Recent work focuses on privacy-preserving techniques that support secure inference without sacrificing performance. Federated learning improves privacy by keeping data local and sharing only model updates, but risks remain as updates can still reveal sensitive information. [191] proposed a homomorphic encryption framework to compute model updates securely. Similarly, [192] introduced PFMLP, using partially homomorphic encryption during federated learning with minimal accuracy impact. In 2023, [193] proposed a framework combining homomorphic encryption and random privacy masks to protect raw input data with low overhead. [194] also integrated secure multiparty computation (SMC) with homomorphic encryption to enhance privacy in federated systems.
Another concern is membership inference attacks, where models trained on private data collections can leak user information [195]. [195] proposed encrypting model updates using homomorphic encryption to prevent such leakage. [196] developed a multi-key encryption scheme ensuring updates remain inaccessible to any single participant. Data ownership and control are also challenges—once data is uploaded to the cloud, users often lose control. [197] addressed this by decentralizing client data in IoT settings and defending against collusion attacks. Privacy regulations like GDPR further complicate data handling. [198] proposed a framework combining differential privacy and homomorphic encryption to ensure GDPR compliance while maintaining model utility.
@@ V-H Communication Model for Edge VLMs
How does one enable communication resources efficient edge VLMs? Training requires significant communication resources, especially for distributed VLMs at the network edge. Offloading model processing to the cloud adds communication overhead. [199] proposed a distributed CNN framework to reduce memory usage and communication in edge-cloud setups. [200] introduced a framework to optimize data flow between edge and cloud. Energy consumption is another key challenge. In 2023, [201] proposed an energy-efficient NLP framework using heterogeneous memory to lower power use while maintaining performance. [202] developed EdgePipe, a distributed framework using pipeline parallelism to improve inference efficiency, achieving speedups without accuracy loss.
Real-time inference is critical for edge applications. [203] proposed DeViT, which decomposes large vision transformers into smaller models for collaborative inference on edge devices, reducing latency and communication overhead. [202] showed that pipeline parallelism can also boost inference on heterogeneous devices with minimal accuracy drop. Generalizing VLMs across diverse hardware is another challenge. [204] proposed DCA-NAS, enabling NAS for various hardware configurations, allowing tailored model designs. [202] also demonstrated pipeline parallelism’s adaptability to heterogeneous systems, improving performance and flexibility.